Hi this is a student of class 12. Can I plz get the ozonolysis reaction.? Wuth example

Grade:12

1 Answers

Arun
5 years ago

If you observe the reaction of pattern of the alkenes, say H2C=CH2, you will notice, in most of the reaction the second bond or the pi-bond breaks. For example, if H2C=CH2 reacts with water (H-OH) it creates alcohol, H3C-C(OH)H2.

In ozonolysis, not only pi-bond is broken, but also the sigma bond breaks in presence of ozone. This is called oxidative cleavage.

H2C=CH2 gets broken in two radicals H2C= and =CH2. The oxygen atoms gets connected with the C atoms formting C=O i.e. the ketone functional group.

H2C=CH2 + O3 -> 2 H2C=O

This reactions are done using reducing any reducing agent line Zinc. The third O atom of the Ozone gets connected with the Zinc.

If the C atoms are connected with any R- group, rather than only H atoms, it can also form aldehydes when reacting with Ozone.

For example,

H-C(CH3) = CH2 + O3 -> H-C(CH3)= + =CH2 + O3 -> (H3C)-CHO + O=CH2